OpenStudy (anonymous):

Solve the following system by solving one of the equations for x or y and then using the substitution method. 5x+y=4 x-3y=4 @geerky42

OpenStudy (anonymous):

These questions are so funnnnnnnn

OpenStudy (anonymous):

Glad YOU think so lol you can help me do them then lol

OpenStudy (anonymous):

Wish it wasn't so late in the night, where I'm too lazy, but I really want to do this. SO, just isolate x in the second equation since it seems easier, or quicker.

OpenStudy (anonymous):

x=3y+4

OpenStudy (anonymous):

Now take that and substitute it into the first equation, do you know what I mean?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

5(3y+4)=4

OpenStudy (anonymous):

You missed something, do you know what?

OpenStudy (anonymous):

+y

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Now, rewrite it again ^.^

OpenStudy (anonymous):

Thank you cuz that just eff'ed up my whole answer! lmaoo 5(3y+4)+y=4 15y+20+y=4 15y+y=-16 16y=-16 y=-1?

OpenStudy (anonymous):

yay now sub that in one of the equations and solve for x

OpenStudy (anonymous):

Then you'll have x and y, and you'll be done :D

OpenStudy (anonymous):

x-3(-1)=4 Will the 3 turn positive?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

x=1! 1 & -1

OpenStudy (anonymous):

Yay, pretty simple right?

OpenStudy (anonymous):

ehhh, sometimes lol